A person drives 390 miles on a stretch of road. Half the distance is driven traveling 5 miles per hour below the speed limit, and half the distance is driven traveling 5 miles per hour above the speed limit. If the time spent traveling at the slower speed exceeds the time spent traveling at the faster speed by 24 minutes, find the speed limit.
Answer by mananth(16949)

first part x-5 mph 195 miles
second part x+5 mph 195 miles
195/(x-5)-195/(x+5) = 0.4 hours
LCD = (x+5)(x-5)
Multiply equation by LCD
195/(x+5)-195/(x-5)= (x+5)(x-5)*0.4
195x+975-195x+975=0.4(x^2-25)
1950=0.4x^2-10
1960=0.4x^2
1960/0.4=x^2
4900 = x^2
x= 70 mph the speed limit